# Note on Min-$k$-Planar Drawings of Graphs

### v4

@@ 1 Introduction
One of the most popular current research directions in graph drawing is going “beyond planarity” [DBLP:journals/csur/DidimoLM19, DBLP:books/sp/20/HT2020]. This somewhat broad direction can be described as considering drawings of graphs in which edges may cross, but the overall pattern of edge crossings is restricted, usually in a local setting. Some of the earliest examples are $1$-planar graphs (every edge may have at most $1$ crossing), and popular extensions nowadays include many families among which we, for example, mention $k$-planar- and fan-planar graphs, or $k$-quasiplanar graphs. These diverse classes often share some nice properties of planar graphs, such as having few edges (e.g., for $1$-planar graphs [https://doi.org/10.1002/mana.19861250122], for $k$-planar graphs [DBLP:journals/combinatorica/PachT97], and for $3$-quasiplanar graphs [DBLP:journals/combinatorica/AgarwalAPPS97]). However, they may differ greatly from planar graphs in other respects; for instance, recognizing $1$-planar graphs is NP-complete [DBLP:journals/algorithmica/GrigorievB07, DBLP:conf/gd/KorzhikM08].
@@
We refer to Section 2 for a precise definition of a drawing of a graph, and of simple and min-$k$-planar drawings. Briefly, by a drawing of a graph (here exclusively in the plane) we mean a topological representation in which edges (as arcs) join their end vertices (as points) and avoid passing through other vertices. Furthermore, every two distinct edges intersect finitely many times, and at most two edges intersect in one point except when it is their common end vertex. In a simple drawing, we additionally require that every two distinct edges intersect in at most one point – a crossing or a common end.
@@
It is very common that research papers assume only simple drawings, for convenience, but it is sometimes not quite clear whether this assumption is made “without loss of generality,” or whether it is a significant restriction on the kinds of drawings considered. For instance, when studying the crossing number of a graph (the minimum total number of crossings over all drawings), one can quite easily restrict attention to simple drawings without loss of generality, but such restriction is not possible when studying the odd crossing number [DBLP:journals/dcg/PelsmajerSS08]. On a different note, so-called fan-planar graphs can be assumed to have a simple fan-planar drawing without loss of generality, but the proof [DBLP:journals/jgaa/KlemzKRS23] is highly nontrivial.
@@
Consider $k$-planar graphs, which are graphs admitting a drawing in which no edge carries more than $k$ crossings. (The same concept is also known as the local crossing number [schaefer2017crossing] or the crossing parameter [DBLP:journals/algorithmica/GrigorievB07].) The seminal paper of Pach and Tóth [DBLP:journals/combinatorica/PachT97] explicitly requires simple $k$-planar drawings, while, e.g., “algorithmic” papers Grigoriev and Bodlaender [DBLP:journals/algorithmica/GrigorievB07] and Korzhik and Mohar [DBLP:conf/gd/KorzhikM08] deal with general $k$-planar drawings (in fact, [DBLP:conf/gd/KorzhikM08] mentions that any $1$-planar drawing can be turned into a simple one) and Ackerman [DBLP:journals/comgeo/Ackerman19] distinguishes the cases. A recent survey on beyond planarity [DBLP:journals/csur/DidimoLM19] unfortunately does not explicitly address this issue, and it mostly only implicitly restricts to results about simple drawings in this respect.
@@
To illustrate the potential problem points (of unwary mixing general and simple $k$-planar drawings in research), we mention, e.g., [DBLP:conf/gd/Bekos0R16] which formulates results about general $3$-planar drawings, but importantly uses a lemma of [DBLP:journals/combinatorica/PachT97] which, in unmentioned fact, relies on the assumption of a simple drawing. In this particular case of [DBLP:conf/gd/Bekos0R16], as well as in other papers which deal with $k$-planar drawings for only $k\leq 3$, there is no reall problem since every general $k$-planar graph for $k\leq 3$ admits a simple $k$-planar drawing, as shown already by Pach et al. [DBLP:journals/dcg/PachRTT06]. On the other hand, for every $k\geq 4$ there exist $k$-planar graphs which have no simple $k$-planar drawing, e.g., Schaefer [schaefer2017crossing, Chapter 7].
@@
Concerning the new related concept of min-$k$-planar graphs, which are graphs admitting a drawing in which every pair of crossing edges has one of the two edges with at most $k$ crossings; the introductory paper by Binucci et al. [DBLP:conf/gd/BinucciBBDDHKLMT23, DBLP:journals/corr/abs-2308-13401] requires simple drawings by the definition. However, min-$k$-planar drawings may also be understood in the general (non-simple) setting, and the difference between the general and the simple settings is much more striking than in the case of $k$-planar graphs.
@@
Namely, we prove (Theorem 2.1) that for arbitrarily large fixed $k$ there exist graphs that are min-$2$-planar without restricting to simple drawings, but which have no simple min-$k$-planar drawing. Alternatively, counterexample graphs with a min-$3$-planar drawing in which no two edges sharing a common vertex cross can also be constructed. In other words, the concepts of simple and general min-$k$-planar drawings always significantly differ, except in the trivial case of $k=1$ (in which we can easily simplify any min-$1$-planar drawing, cf. Section 2), and they differ for $k\geq 3$ even if we forbid general drawings in which two edges with a common vertex cross.
@@
In the course of proving this result, we develop a technical tool (Section 3) which suitably constrains possible min-$k$-planar drawings of graphs within a rigid “frame”, and we suggest this tool can be useful in further research of the properties of min-$k$-planar graphs.
@@ 2 Min-$k$-planar Drawings and Graphs
For further concepts, the definition of a ‘crossing’ needs a careful clarification. If $k>2$ edges together cross in one point, then this situation, strictly, counts as ($k\choose 2$ crossings of pairs of these edges (see, e.g., [DBLP:journals/siamcomp/CabelloM13]). Since a slight local perturbation of the involved edges avoids common crossings of triples without adding any new crossing, for our purposes, we may simply discard the possibility that $k>2$ edges together cross in one point, which is a common approach in literature. Therefore, without loss of generality, we require that, in every drawing, no three edges cross in the same point, unless stated otherwise.
In a simple drawing $\mathcal{G}$ (otherwise known as a good drawing [cd98262b-e086-3d8d-b726-cd66f05da520, schaefer2017crossing], but we wish to adhere to the recent terminology), crossings are allowed, but (again) no three edges cross in the same point, no two edges have more than one crossing in common, and no two adjacent edges (i.e., with a common end vertex) cross. Hence, in our setting, a simple drawing $\mathcal{G}$ is defined such that, for every $e\not=f\in E(G)$, we have $|\mathcal{G}(e)\cap\mathcal{G}(f)|\leq 1$, except that $|\mathcal{G}(e)\cap\mathcal{G}(f)|=2$ when $e$ and $f$ are parallel edges (the latter is irrelevant for simple graphs).
@@
To complement Theorem 2.1, we resolve the remaining trivial cases in Section 2. The statement is simply illustrated in (a) Figure 1 and (b) Figure 2, but one has to be a bit careful with covering all possible cases which may occur. Full technical details are left for the Appendix.
{proposition2rep}
@@
Figure 1: An illustration of Section 2 a); up to symmetry between $e$ and $f$ (with common end vertex $x$), the edge $e$ carries no other crossing than the point $y$, and so one can draw an uncrossed arc $e^{\prime}_{1}$ tightly along the segment $e_{1}\subseteq e$ from $x$ to $y$. When redrawing from $e$ to $e^{\prime\prime}$ and from $f$ to $f^{\prime\prime}$ (using $e^{\prime}_{1}$), the crossing at $y$ is eliminated and no new crossings are added to any edge in the picture.
@@
Figure 2: An illustration of Section 2 b); now we have two crossings $x$ and $y$ of the same pair $e$ and $f$ of edges, and there are no more crossings on $e$. Similarly to Figure 1, when redrawing from $e$ to $e^{\prime\prime}$ and from $f$ to $f^{\prime\prime}$ (using the uncrossed arc $e^{\prime}_{1}$), at least one of the crossings at $x$ or $y$ is eliminated and no new crossings are added to any edge in the picture.
Proof 2.2.
@@
Proof 2.4.
We define the anchored graph $(G_{k},A_{k})$ as depicted in Figure 3. $G_{k}$ is a disjoint union of two induced matchings $M_{1}$ and $M_{2}$ of $k+1$ edges each, and of two induced stars $S_{1}$ and $S_{2}$, where $S_{1}$ has centre $a_{2}$ and $k+1$ leaves including $a_{1},b_{1}$ and $S_{2}$ has centre $c_{2}$ and $k+2$ leaves in the set $C_{3}\cup\{c_{1}\}$ where $|C_{3}|=k+1$. The anchor set is $A_{k}=V(G_{k})\setminus\{c_{2}\}$ ordered as in Figure 3. Figure 3 also shows an anchored min-$2$-planar drawing of $(G_{k},A_{k})$ (however, $a_{1}a_{2}$ crosses $b_{1}a_{2}$ there).
@@
Figure 3: An anchored min-$2$-planar drawing of the anchored graph $(G_{k},A_{k})$ (here for $k=3$, with the heavy edges drawn thick) as used in the proof of Lemma 2.3. With general $k\geq 2$, at least $k-1$ of the edges of the star $S_{3}$ with center $a_{2}$ follow the depicted route of the red edge $a_{2}b_{1}$.
@@
Figure 4: An anchored general min-$3$-planar drawing of the anchored graph $(G^{\prime}_{k},A^{\prime}_{k})$ (here for $k=4$, with the heavy edges drawn thick) as used in the proof of Lemma 2.3. For general $k\geq 3$, at least $k-2$ of the edges of the matching $M_{3}$ follow the route of the red edge $b_{1}b_{2}$.
@@ 3 Technical Proofs
Proof 3.2 (Proof sketch).
Let the length-$2$ paths (with ends in $V(G)$ ) from the definition of the $t$-amplification $G^{t}$ be called double edges of $G^{t}$. Observe that if, for some two edges $e,f\in E(G)$, sufficiently many of the double edges coming from $e$ cross many of those coming from $f$, then $\mathcal{G}$ is not min-$k$-planar. Using a Ramsey-type argument, we can thus conclude that large amounts of double edges coming from $e$ and from $f$ are mutually noncrossing, and this finding can be iteratively extended to all pairs of edges of $G$ at once. This gives the desired planar subdrawing $\mathcal{G}^{\prime}\subseteq\mathcal{G}$. Full technical details are left for the Appendix.
{lemma2rep}
@@
Figure 5: An illustration of the graph $H_{2}$ from the proof of Section 3. The three black circled vertices are the designated anchors $A$ (here $|A|=3$). The subgraph $H_{2}^{\prime}$ is in thick black and $H_{2}^{\prime\prime}$ in magenta colour. All magenta edges get $t$-amplified in the construction of $H$, and consequently, the black edges forced to cross them will be heavy in any min-$k$-planar drawing of $H$ by Section 3.
Proof 3.3.
@@
Furthermore, let $H_{2}^{\prime}\subseteq H_{2}$ be the subgraph formed by the original edges of $H_{0}$, and $H_{2}^{\prime\prime}=H_{2}\setminus E(H_{2}^{\prime})$. Observe one important property of planar $H_{2}^{\prime\prime}$; for every planar drawing of $H_{2}^{\prime\prime}$ and each edge $uv\in E(H_{2}^{\prime})$, the points $u$ and $v$ are separated by a cycle (e.g., the dual cycle of $u\not=w_{2}$) in $H_{2}^{\prime\prime}$. Let now $t=t(H_{2}^{\prime\prime},k,k+1)$ be as in Section 3. We construct $H:=H_{2}^{\prime}\cup H_{2}^{t}$ where $H_{2}^{t}$ is the $t$-amplification of $H_{2}^{\prime\prime}$. Obviously, $(H,A)$ has an anchored simple min-$1$-planar drawing, e.g., one following Figure 5.
Consider now a min-$k$-planar drawing $\mathcal{H}$ of $H\cup G$, and denote by $\mathcal{H}_{0}\subseteq\mathcal{H}$ the subdrawing of $H_{2}^{t}$. By Section 3, $\mathcal{H}_{0}$ contains a planar subdrawing of a $(k+1)$-amplification of $H_{2}^{\prime\prime}$. By the mentioned property of $H_{2}^{\prime\prime}$, every edge of $H_{2}^{\prime}$ thus has to cross at least $k+1$ edges of $H_{2}^{t}$ and is heavy. In particular, the subdrawing of $H_{2}^{\prime}$ within $\mathcal{H}$ is planar (and so has to look as in Figure 5).
@@
Proof 3.4 (Proof sketch).
We begin with a graph $H_{0}$ being a sufficently large double wheel, i.e., a graph made from a cycle by adding two central vertices $w_{1},w_{2}$ adjacent to all cycle vertices. Let $H_{0}^{*}$ be the planar dual of $H_{0}$, and $H_{1}$ be constructed from $H_{0}\cup H_{0}^{*}$ by adding extra edges $uv\in E(H_{1})$ for every pair $u\in V(H_{0})$ and $v\in V(H_{0}^{*})$ such that $u$ is a vertex incident to the face of $H_{0}$ represented by $v$. We construct $H_{2}$ from $H_{1}$ by splitting the central vertex $w_{2}$ into $a$ vertices incident to equally split consecutive spoke edges of $H_{0}$, as sketched in Figure 5. The vertices split from $w_{2}$, in their clockwise order, define the anchor set $A$.
Let $H_{2}^{\prime}\subseteq H_{2}$ be the subgraph formed by the original edges of $H_{0}$, and $H_{2}^{\prime\prime}=H_{2}\setminus E(H_{2}^{\prime})$. For sufficiently large $t$, we construct $H:=H_{2}^{\prime}\cup H_{2}^{t}$ where $H_{2}^{t}$ is the $t$-amplification of $H_{2}^{\prime\prime}$. Obviously, $(H,A)$ has an anchored simple min-$1$-planar drawing, e.g., one following Figure 5.
Now, in any min-$k$-planar drawing $\mathcal{H}$ of $H\cup G$, by Section 3, there is a planar subdrawing of a $(k+1)$-amplification of $H_{2}^{\prime\prime}$ which, hence, crosses every edge of $H_{2}^{\prime}$ in $\mathcal{H}$ at least $(k+1)$-times. From this fact, and assuming sufficiently dense $H_{2}^{\prime}$, we derive that the central vertex $w_{1}$ of $H_{2}^{\prime}$ is connected to every anchor vertex in $A$ by a path which is uncrossed in $\mathcal{H}$. By “inverting” a local neighbourhood of $w_{1}$ into the outer face, we thus get an anchored subdrawing of $(G,A)$ or its mirror image from this. Full details are left for the Appendix.
@@
We take the anchored graph $(G_{k},A_{k})$, $k\geq 2$, and the anchored graph $(G^{\prime}_{k},A^{\prime}_{k})$, $k\geq 3$, from Lemma 2.3 for (a) and (b), respectively, and plug them into Section 3 as $G$. The rest is an immediate consequence of the previous statements.

### v5

@@ 1 Introduction
One of the most popular current research directions in graph drawing is going “beyond planarity” [7, 10]. This somewhat broad direction can be described as considering drawings of graphs in which edges may cross, but the overall pattern of edge crossings is restricted, usually in a local setting. Some of the earliest examples are $1$-planar graphs (every edge may have at most $1$ crossing), and popular extensions nowadays include many families among which we, for example, mention $k$-planar- and fan-planar graphs, or $k$-quasiplanar graphs. These diverse classes often share some nice properties of planar graphs, such as having few edges (e.g., for $1$-planar graphs [18], for $k$-planar graphs [14], and for $3$-quasiplanar graphs [2]). However, they may differ greatly from planar graphs in other respects; for instance, recognizing $1$-planar graphs is NP-complete [9, 12].
@@
We refer to Section 2 for a precise definition of a drawing of a graph, and of simple and min-$k$-planar drawings. Briefly, by a drawing of a graph (here exclusively in the plane) we mean a topological representation in which edges (as arcs) join their end vertices (as points) and avoid passing through other vertices. Furthermore, every two distinct edges intersect finitely many times, and at most two edges intersect in one point except when it is their common end vertex. In a simple drawing, we additionally require that every two distinct edges intersect in at most one point – a crossing or a common end. (However, one has to be careful with this simplified definition when considering a simple drawing of a non-simple graph; then two parallel edges share two common ends.)
@@
It is very common that research papers assume only simple drawings, for convenience, but it is sometimes not quite clear whether this assumption is made “without loss of generality,” or whether it is a significant restriction on the kinds of drawings considered. For instance, when studying the crossing number of a graph (the minimum total number of crossings over all drawings), one can quite easily restrict attention to simple drawings without loss of generality, but such restriction is not possible when studying the odd crossing number [16]. On a different note, so-called fan-planar graphs can be assumed to have a simple fan-planar drawing without loss of generality, but the proof [11] is highly nontrivial.
@@
Consider $k$-planar graphs, which are graphs admitting a drawing in which no edge carries more than $k$ crossings. (The same concept is also known as the local crossing number [17] or the crossing parameter [9].) The seminal paper of Pach and Tóth [14] explicitly requires simple $k$-planar drawings, while, e.g., “algorithmic” papers Grigoriev and Bodlaender [9] and Korzhik and Mohar [12] deal with general $k$-planar drawings (in fact, [12] mentions that any $1$-planar drawing can be turned into a simple one) and Ackerman [1] distinguishes the cases. A recent survey on beyond planarity [7] unfortunately does not explicitly address this issue, and it mostly only implicitly restricts to results about simple drawings in this respect.
@@
To illustrate the potential problem points (of unwary mixing general and simple $k$-planar drawings in research), we mention, e.g., [3] which formulates results about general $3$-planar drawings, but importantly uses a lemma of [14] which, in unmentioned fact, relies on the assumption of a simple drawing. In this particular case of [3], as well as in other papers which deal with $k$-planar drawings for only $k\leq 3$, there is no reall problem since every general $k$-planar graph for $k\leq 3$ admits a simple $k$-planar drawing, as shown already by Pach et al. [13]. On the other hand, for every $k\geq 4$ there exist $k$-planar graphs which have no simple $k$-planar drawing, e.g., Schaefer [17, Chapter 7].
@@
Concerning the new related concept of min-$k$-planar graphs, which are graphs admitting a drawing in which every pair of crossing edges has one of the two edges with at most $k$ crossings; the introductory paper by Binucci et al. [4, 5] requires simple drawings by the definition. However, min-$k$-planar drawings may also be understood in the general (non-simple) setting, and the difference between the general and the simple settings is much more striking than in the case of $k$-planar graphs.
@@
Namely, we prove (Theorem 2.1) that for arbitrarily large fixed $k$ there exist graphs that are min-$2$-planar without restricting to simple drawings, but which have no simple min-$k$-planar drawing. Alternatively, counterexample graphs with a min-$3$-planar drawing in which no two edges sharing a common vertex cross can also be constructed. In other words, the concepts of simple and general min-$k$-planar drawings always significantly differ, except in the trivial case of $k=1$ (in which we can easily simplify any min-$1$-planar drawing, cf. Proposition 2.2), and they differ for $k\geq 3$ even if we forbid general drawings in which two edges with a common vertex cross.
@@
In the course of proving this result, we develop a technical tool (Lemma 3.3) which suitably constrains possible min-$k$-planar drawings of graphs within a rigid “frame”, and we suggest this tool can be useful in further research of the properties of min-$k$-planar graphs.
@@ 2 Min-$k$-planar Drawings and Graphs
For further concepts, the definition of a ‘crossing’ needs a careful clarification. If $k>2$ edges together cross in one point, then this situation, strictly, counts as ($k\choose 2$ crossings of pairs of these edges (see, e.g., [6]). Since a slight local perturbation of the involved edges avoids common crossings of triples without adding any new crossing, for our purposes, we may simply discard the possibility that $k>2$ edges together cross in one point, which is a common approach in literature. Therefore, without loss of generality, we require that, in every drawing, no three edges cross in the same point, unless stated otherwise.
One may, likewise, deal with possible tangential crossings in a drawing, which can be discarded by a slight local perturbation as well, and so exclude tangential crossings by the definition (as many papers do for convenience), but here we stay on the more general side and do not exclude them.
In a simple drawing $\mathcal{G}$ (otherwise known as a good drawing [8, 17], but we wish to adhere to the recent terminology), crossings are allowed, but (again) no three edges cross in the same point, no two edges have more than one crossing in common, and no two adjacent edges (i.e., with a common end vertex) cross. Hence, in our setting, a simple drawing $\mathcal{G}$ is defined such that, for every $e\not=f\in E(G)$, we have $|\mathcal{G}(e)\cap\mathcal{G}(f)|\leq 1$, except that $|\mathcal{G}(e)\cap\mathcal{G}(f)|=2$ when $e$ and $f$ are parallel edges (the latter is irrelevant for simple graphs).
@@
To complement Theorem 2.1, we resolve the remaining trivial cases in Proposition 2.2.
Proposition 2.2.
@@
Figure 1: An illustration of Proposition 2.2 a); up to symmetry between $e$ and $f$ (with common end vertex $x$), the edge $e$ carries no other crossing than the point $y$, and so one can draw an uncrossed arc $e^{\prime}_{1}$ tightly along the segment $e_{1}\subseteq e$ from $x$ to $y$. When redrawing from $e$ to $e^{\prime\prime}$ and from $f$ to $f^{\prime\prime}$ (using $e^{\prime}_{1}$), the crossing at $y$ is eliminated and no new crossings are added to any edge in the picture.
@@
Figure 2: An illustration of Proposition 2.2 b); now we have two crossings $x$ and $y$ of the same pair $e$ and $f$ of edges, and there are no more crossings on $e$. Similarly to Figure 1, when redrawing from $e$ to $e^{\prime\prime}$ and from $f$ to $f^{\prime\prime}$ (using the uncrossed arc $e^{\prime}_{1}$), at least one of the crossings at $x$ or $y$ is eliminated and no new crossings are added to any edge in the picture.
Proof 2.3.
@@
Proof 2.5.
We define the anchored graph $(G_{k},A_{k})$ as depicted in Figure 3. $G_{k}$ is a disjoint union of two induced matchings $M_{1}$ and $M_{2}$ of $k+1$ edges each, and of two induced stars $S_{1}$ and $S_{2}$, where $S_{1}$ has center $a_{2}$ and $k+1$ leaves including $a_{1},b_{1}$ and $S_{2}$ has center $c_{2}$ and $k+2$ leaves in the set $C_{3}\cup\{c_{1}\}$ where $|C_{3}|=k+1$. The anchor set is $A_{k}=V(G_{k})\setminus\{c_{2}\}$ ordered as in Figure 3. Figure 3 also shows an anchored min-$2$-planar drawing of $(G_{k},A_{k})$ (however, $a_{1}a_{2}$ crosses $b_{1}a_{2}$ there).
@@
Figure 3: An anchored general min-$2$-planar drawing of the anchored graph $(G_{k},A_{k})$ (here for $k=3$, with the heavy edges drawn thick) as used in the proof of Lemma 2.4. With general $k\geq 2$, at least $k-1$ of the edges of the star $S_{1}$ with center $a_{2}$ follow the depicted route of the red edge $a_{2}b_{1}$, while crossing the edge $a_{2}a_{1}$.
@@
Figure 4: An anchored general min-$3$-planar drawing of the anchored graph $(G^{\prime}_{k},A^{\prime}_{k})$ (here for $k=4$, with the heavy edges drawn thick) as used in the proof of Lemma 2.4. For general $k\geq 3$, at least $k-2$ of the edges of the matching $M_{3}$ follow the route of the red edge $b_{1}b_{2}$. These edges cross the heavy edge $a_{1}a_{2}$ twice, and so the drawing is neither simple nor general min-$2$-planar.
@@ 3 Technical Proofs
Lemma 3.3.
@@
Figure 5: An illustration of the graph $H_{2}$ from the proof of Lemma 3.3. The three black circled vertices are the designated anchors $A$ (here $A=\{w_{2}^{1},w_{2}^{2},w_{2}^{3}\}$). The subgraph $H_{2}^{\prime}$ is in thick black and $H_{2}^{\prime\prime}$ in magenta colour. All magenta edges get $t$-amplified in the construction of $H$, and consequently, the black edges forced to cross them will be heavy in any min-$k$-planar drawing of $H$ by Lemma 3.1.
Proof 3.4.
@@
Furthermore, let $H_{2}^{\prime}\subseteq H_{2}$ be the subgraph formed by the original edges of $H_{0}$, and $H_{2}^{\prime\prime}=H_{2}\setminus E(H_{2}^{\prime})$. Observe one important property of planar $H_{2}^{\prime\prime}$; for every planar drawing of $H_{2}^{\prime\prime}$ and each edge $uv\in E(H_{2}^{\prime})$, the points $u$ and $v$ are separated by a cycle (e.g., the dual cycle of $u\not=w_{2}$) in $H_{2}^{\prime\prime}$. Let now $t=t(H_{2}^{\prime\prime},k,k+1)$ be as in Lemma 3.1. We construct $H:=H_{2}^{\prime}\cup H_{2}^{t}$ where $H_{2}^{t}$ is the $t$-amplification of $H_{2}^{\prime\prime}$. Obviously, $(H,A)$ has an anchored simple min-$1$-planar drawing, e.g., one following Figure 5.
Consider now a min-$k$-planar drawing $\mathcal{H}$ of $H\cup G$, and denote by $\mathcal{H}_{0}\subseteq\mathcal{H}$ the subdrawing of $H_{2}^{t}$. By Lemma 3.1, $\mathcal{H}_{0}$ contains a planar subdrawing of a $(k+1)$-amplification of $H_{2}^{\prime\prime}$. By the mentioned property of $H_{2}^{\prime\prime}$, every edge of $H_{2}^{\prime}$ thus has to cross at least $k+1$ edges of $H_{2}^{t}$ and is heavy. In particular, the subdrawing of $H_{2}^{\prime}$ within $\mathcal{H}$ is planar (and so has to look as in Figure 5).
@@
We take the anchored graph $(G_{k},A_{k})$, $k\geq 2$, and the anchored graph $(G^{\prime}_{k},A^{\prime}_{k})$, $k\geq 3$, from Lemma 2.4 for (a) and (b), respectively, and plug them into Lemma 3.3 as $G$. The rest is an immediate consequence of the previous statements.